OpenStudy (mrhoola):

Can any bdy solve the last part of this one : Diff. Eq , separable variable method... dP/dt= P-P^2 * dont forget to consider the chain rule when you integrate *

OpenStudy (unklerhaukus):

\[\frac{\text dP}{\text dt}= P-P^2\] \[\frac{\text dP}{P-P^2}= \text dt\] \[\int\frac{\text dP}{P-P^2}= \int\text dt\]

OpenStudy (callisto):

For the left: \[\int \frac{dP}{P-P^2} = \int \frac{dP}{P(1-P)}=\int \frac{1}{P}+\frac{1}{1-P}dP\]